# Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models

### v1

@@ Abstract
LLMs have demonstrated impressive proficiency in generating coherent and high-quality text, making them valuable across a range of text-generation tasks. However, rigorous evaluation of this generated content is crucial, as ensuring its quality remains a significant challenge due to persistent issues such as factual inaccuracies and hallucinations. This paper introduces two fine-tuned general-purpose LLM autoevaluators, REC-12B and REC-70B—specifically designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, and completeness. These models not only provide ratings for these metrics but also offer detailed explanations and verifiable citations, thereby enhancing trust in the content. Moreover, the models support various citation modes, accommodating different requirements for latency and granularity. Extensive evaluations on diverse benchmarks demonstrate that our general-purpose LLM auto-evaluator, REC-70B, outperforms state-of-the-art LLMs, excelling in content evaluation by delivering better quality explanations and citations with minimal bias. It achieves Rank #1 as a generative model on the RewardBench leaderboard [^1] https://huggingface.co/spaces/allenai/reward-bench under the model name TextEval-Llama3.1-70B. Our REC dataset and models are released at https://github.com/adelaidehsu/REC.
@@ 1 Introduction
Figure 1: Illustration of task prompt, context, response, and model outputs of Rating, Explanation, and Citations (REC). Here we use summarization evaluation as an example.
@@
Large Language Models (LLMs) have demonstrated impressive capabilities in generating high quality coherent text and are deployed in applications for various text-generation tasks (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Touvron et al., 2023). In order for LLMs to provide up-to-date information or to perform knowledge-intensive tasks (Lewis et al., 2020), Retrieval Augmented Generation (RAG) system (Chen et al., 2017; Borgeaud et al., 2021; Izacard et al., 2022; Guu et al., 2020) has been used widely. RAG involves first retrieving document chunks that are relevant to the generation task from an external knowledge source, and performing the generation using an knowledge-augmented prompt. However, it’s of paramount importance to ensure that the generated text is reliable and can be trusted by a human as LLMs often suffer from factual incorrectness and hallucination of knowledge (Ji et al., 2022; Zhang et al., 2023; Shuster et al., 2021). Towards this goal, we propose the generated text needs to be evaluated along various dimensions shown below:
  • Faithfulness: Did the LLM generate factually correct response given the context? (Rating + Explanation task)
@@
  • Instruction Following: Did the LLM generate response follow the instructions provided in the prompt? (Rating + Explanation task)
  • Coherence: Did the LLM generate a coherent response? (Rating + Explanation task)
  • Completeness: Did the LLM generate a complete response including all details? (Rating + Explanation task)
@@
  • Citation: If the LLM generated factually correct response, can we provide evidence for where that response came from? (Citation task)
@@
In this paper, we introduce fine-tuned models for the evaluation tasks listed above that can form the basis for all trust-related evaluations for generative AI applications. The model was fine-tuned to not only provide ratings for the metrics but also explanations + citations for why it rated the generation so. For instance, if the faithfulness metric is rated as “factually correct", the model will output explanation with verifiable citations as to why it rated the generation as factually correct. An example of the model output is shown in Figure 1.
The listed metrics are essentially Rating and Explanation tasks for our fine-tuned models which takes as input: (1) another task LLM’s hydrated prompt (2) a generated response by the task LLM. The output would be a rating and an associated explanation for that rating as shown in Figure 2.
Figure 2: Model Inputs and Outputs when providing Rating with Explanations and citations.
The Citation task takes as input (1) a provided context (2) a generated response. The output would be a citation tagged response that provides evidence for parts of generated text as shown in Figure 2.
To differentiate from the content quality citations  [^2] For content quality citation task, we cite from another task LLM’s hydrated prompt (context) and include the citations with the generated explanation (response) by the LLM autoevaluator. we obtain when using our model as an autoevaluator for verifiable evaluation by performing Rating, Explanation and Citation, we emphasize when performing the Citation task alone, our model is able to conduct general RAG citations, where we cite from the retrieved chunks (context) and tag the citations with the generated answer (response). For both kinds, we support various citation modes as listed below (See Appendix A for examples on RAG citation modes):
  • Post-fix citation: If the generation is few sentences long, we place the citation at the end of the response and refer to the context id.
  • Post-fix citation mode with snippet: If the generation is few sentences long, we place the citation at the end of the response and refer to the exact sentence within that context that supports the generation.
  • Inline citation: If the generation is long, we place the citation at appropriate locations within the generation and refer to the context id.
  • Inline citation mode with snippet: If the generation is long, we place the citation at appropriate locations within the generation and refer to the exact sentence within that context that supports the generation.
These different modes of citations were designed to cater to different trade-offs between latency and granularity of citation. For instance, the Post-fix citation mode would be the fastest as there is no need for our fine-tuned model to generate claims or snippet. It simply has to generate the reference to the cited context. On the other hand, the inline citation mode with snippet has to place the citation inline with the generated response and also point to a snippet within the corresponding cited context. This mode is the most granular but can also increase latency as it needs to generate more output tokens.
@@
Our model is the first to enable citation in both automatic evaluation and general task output with scalability and efficiency, while prior work often solely focus on providing rating and explanation in automatic evaluation (Zhu et al., 2023; Li et al., 2023a; Wang et al., 2023b; Vu et al., 2024a), or requires iterative prompting (Sun et al., 2023; Cao and Wang, 2024; Ye et al., 2023) or human annotation (Malaviya et al., 2024) for general citations generation. Our contributions include the following:
@@
  • A novel general-purpose LLM autoevaluator that comes in two sizes: 12B and 70B, that can generate better quality Rating, Explanation and Citations (REC), with little to no trade-off in general instruction task performance evaluated on various public benchmark datasets and our own dataset described below.
@@
  • A curated dataset for citations and explanations fine-tuning to facilitate future research, which is the first public dataset containing both content quality citations and RAG citations.
@@
  • A single model that can perform different modes of citation to cover the trade-off between latency and granularity of citation. Such model with generalized capability largely simplify the deployment complexity in production.
@@
We provide extensive evaluations of our REC models on diverse benchmarks evaluating LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); (4) cognitive bias: CoBBLEr (Koo et al., 2024), against 7 state-of-the-art (SOTA) LLMs, such as GPTs and Claude. We show our REC models outperforms the state-of-the-art baselines on these benchmarks, and that the baselines often struggle to provide accurate citations, undermining their reliability. Specifically, our REC-70B model achieves Rank #1 as a generative model on the RewardBench leaderboard under the model name TextEval-Llama3.1-70B. The model outperforms the second ranked model in three dimensions: Chat Hard, Safety, Reasoning, which are key capabilities required to generate accurate explanation and citation outputs.
@@ 2 Related Work
Providing citations to attributable information source (Liu et al., 2023a; Gao et al., 2023) has been proposed as one of the approaches to mitigate hallucination and improve the factuality of LLM generation. Such work falls into the general RAG citation setting defined in our work, where the aim is to generate verifiable content through citing from the provided document chunks. Initial efforts fine-tune LLMs using either human-written examples (Nakano et al., 2021) or machine-generated examples verified by human (Menick et al., 2022), but their privately maintained training data limits further research. With the advent of more capable LLMs (OpenAI, 2023; Jiang et al., 2023), most existing work relies on zero-shot prompting or few-shot prompting to cite document chunks during generation (Kamalloo et al., 2023; Gao et al., 2023; Liu et al., 2023a), although the quality of their generated citations leaves large room for improvement (Malaviya et al., 2024). Other work utilizes an additional natural language inference (NLI) model to add citations (Gao et al., 2022; Chen et al., 2023). To improve upon prior work, we propose a fine-tuned LLM that can generate better-grounded responses supported with citations of various modes, with the fine-tuning dataset released to the public to facilitate future research and no additional NLI models needed.
@@
Collecting human annotations to evaluate LLM is not only costly, time-consuming, but hard to replicate (Ouyang et al., 2022; Zheng et al., 2023; Chiang and Lee, 2023), as a result LLMs become a natural automated proxy to evaluate LLM capabilities on various benchmarks (Bai et al., 2024; Bubeck et al., 2023; Chiang et al., 2023; Fu et al., 2023; Liu et al., 2023b; Wang et al., 2023a). Existing LLM autoevaluators often judge LLM outputs by expressing “preference" over outputs from a reference model (Dubois et al., 2024; Li et al., 2023b; Yuchen Lin et al., 2024), or by providing rating and explanation according to a user-defined metric as a direct assessment (Li et al., 2023a; Wang et al., 2023b, 2024b). Closer to our work, Jiang et al. (2024) and Xu et al. (2023) fine-tunes LLMs to generate rating, explanation, and a detailed analysis to pinpoint errors in the response evaluated. Unlike prior work, our fine-tuned model is a more generalizable autoevaluator since it provides content quality citations for both good and bad evaluated responses, in addition to the rating and explanation. This allows users to not only be able to diagnose where errors are from but provide supporting evidence when a generation is good, which are both important to establish trust in the model.
@@ 3 Data Collection
To fine-tune our general-purpose LLM autoevaluator, we meticulously collect a multitask mixture of data encompassing a broad spectrum of LLM capabilities (Section 3.1), and we denote our collected dataset as REC-Data. Due to a lack of publicly available content quality citation datasets, we leverage synthetic data generation using LLM to curate such data (Section 3.2) to facilitate the research community. We perform rule-based automatic quality check followed by a unified task formatting to post-process our data for instruction fine-tuning (Section 3.3).
@@ 3.1 Task Types
To enhance the explainability in the automatic evaluation of our LLM autoevaluator, while maintaining its general instruction-following capability, we gather datasets from a diverse range of task types (See detailed REC-Data distribution in Appendix B). These include:
@@
  • Citation: Evaluate a response alongside the context, and provide citations for verifiable answer attribution.
@@ 3.2 Synthetic data generation
We leverage synthetic data generation for some of the tasks, including pointwise evaluation and citation. For synthetic pointwise evaluation data, we leverage Mistral-7B to rate a response according to one of the 4 metrics listed in Section 1. For synthetic citation data generation, due to the challenging nature of citation tasks, we leverage a more powerful instruction-tuned LLM based on Llama-3.1-70B for both RAG and content quality citations. Example prompts used to generate the synthetic data is provide in Appendix C.
@@ 3.3 Post-processing and Unified Task format
Knowing that LLM-generated outputs is susceptible to hallucination, we perform automatic rule-based content quality check to ensure the quality of the synthetic data. During post-processing, we filter invalid synthetic LLM outputs that have either invalid output formats, or non-factual content cited by checking exact-match of the citation with the context. Our final REC-Data has in total around 140k datapoints, each containing a task prompt and a completion.
ASQA QAMPARI ELI5 AVG
Fluency Correct Citation F1 Correct Citation F1 Fluency Correct Citation F1
Mistral-7B 86.55 19.83 65.31 15.65 16.24 42.43 9.89 47.28 37.90
Mistral-Nemo 80.42 21.49 67.49 17.11 20.85 42.29 15.49 35.61 37.59
Llama-3.1-70B 77.95 25.51 67.07 17.72 16.90 66.07 14.06 36.56 40.23
Claude-3-Opus 89.22 24.08 69.29 0.54 1.92 45.84 11.72 34.69 34.66
GPT-3.5 86.06 23.63 81.02 21.20 19.79 16.35 12.23 51.80 39.01
GPT-4 68.99 25.59 71.73 20.48 25.63 46.98 14.19 43.92 39.69
GPT-4o 79.87 28.73 67.29 20.97 27.11 32.71 17.70 27.11 37.69
REC-12B 81.85 24.36 70.59 18.18 22.71 48.02 16.27 37.60 39.95
REC-70B 79.35 26.06 68.93 18.59 17.48 66.80 16.92 38.79 41.62
Table 1: Fluency, citation quality and correctness on ASQA, QAMPARI, and ELI5 datasets of ALCE

### v2

@@ Abstract
LLMs have demonstrated impressive proficiency in generating coherent and high-quality text, making them valuable across a range of text-generation tasks. However, rigorous evaluation of this generated content is crucial, as ensuring its quality remains a significant challenge due to persistent issues such as factual inaccuracies and hallucination. This paper introduces three fine-tuned general-purpose LLM autoevaluators, REC-8B, REC-12B and REC-70B, specifically designed to evaluate generated text across several dimensions: faithfulness, instruction following, coherence, and completeness. These models not only provide ratings for these metrics but also offer detailed explanation and verifiable citation, thereby enhancing trust in the content. Moreover, the models support various citation modes, accommodating different requirements for latency and granularity. Extensive evaluations on diverse benchmarks demonstrate that our general-purpose LLM auto-evaluator, REC-70B, outperforms state-of-the-art LLMs, excelling in content evaluation by delivering better quality explanation and citation with minimal bias. It achieves Rank #1  [^1] As of Feb 15th, 2025as a generative model on the RewardBench leaderboard [^2] https://huggingface.co/spaces/allenai/reward-bench under the model name TextEval-Llama3.1-70B. Our REC dataset and models are available at https://github.com/adelaidehsu/REC.
@@ 1 Introduction
Figure 1: Illustration of the REC framework for content quality evaluation. The autoevaluator takes in context (a task prompt from another LLM), generation from another LLM, and a user-specified evaluation metric (i.e., completeness). The autoevaluator outputs rating, explanation with citation according to a user-specified citation mode (i.e., inline with context snippet). Citation are extracted verbatim from context as underlined. For details of the alternative citation mode and the evaluation prompt, see Appendix A.
@@
Large Language Models (LLMs) have demonstrated impressive capabilities in generating high quality coherent text and are deployed in applications for various text-generation tasks (Brown et al., 2020; Chowdhery et al., 2022; OpenAI, 2023; Touvron et al., 2023). In order for LLMs to provide up-to-date information or to perform knowledge-intensive tasks (Lewis et al., 2020), Retrieval Augmented Generation (RAG) system (Chen et al., 2017; Borgeaud et al., 2021; Izacard et al., 2022; Guu et al., 2020) has been widely used. RAG involves first retrieving documents that are relevant to the generation task from an external knowledge source, and performing the generation using a knowledge-augmented prompt. However, it is of paramount importance to ensure that the generated text is reliable and can be trusted by a human, as LLMs often suffer from factual inaccuracies and hallucination of knowledge (Ji et al., 2022; Zhang et al., 2023; Shuster et al., 2021). Towards this goal, we propose that the generated text needs to be evaluated along various dimensions shown below:
  • Faithfulness:Is the LLM generated response factually correct given the context?
@@
  • Instruction Following: Does the LLM generated response follow the instructions provided in the prompt?
  • Coherence: Is the LLM generated response coherent?
  • Completeness: Does the LLM generated response include all the necessary details?
@@
  • Citation: If a LLM generated response is factually correct, can we provide evidence for where that response came from?
@@
In this paper, we introduce fine-tuned models for the evaluation tasks listed above that can form the basis for all trust-related evaluations for generative AI applications. The models were fine-tuned to not only provide ratings for the metrics but also explanations + citation for why it rated the generation so. An example of how our models operate for content quality evaluation can be found in Figure 1.
@@
Our models are the first to enable citation in both automatic evaluation (i.e., content quality citation) and general task output (i.e., general RAG citation) with scalability and efficiency, while prior work often focus on providing solely rating and explanation in automatic evaluation (Zhu et al., 2023; Li et al., 2023a; Wang et al., 2023b; Vu et al., 2024a), or require iterative prompting (Sun et al., 2023; Cao and Wang, 2024; Ye et al., 2023) or human annotation (Malaviya et al., 2024) for general citation generation. Our contributions include the following:
@@
  • A novel general-purpose LLM autoevaluator that comes in three sizes: 8B, 12B and 70B. Our LLM autoevaluators can generate improved quality Rating, Explanation and Citation (REC), with little to no trade-off in general instruction task performance.  [^3] All evaluations are limited to English datasets.
@@
  • A curated dataset for citation and explanation fine-tuning to facilitate future research, which is the first public dataset containing both content quality citation  [^4] Citation to support LLM generated explanation for content quality evaluation. See Section 2 for more details. and RAG citation.
@@
  • A single model that can perform different modes of citation to cover the trade-off between latency and granularity of citation.
@@
We provide extensive evaluations of our REC models on diverse benchmarks evaluating LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); (4) cognitive bias: CoBBLEr (Koo et al., 2024), against 7 state-of-the-art (SOTA) LLMs, such as GPTs and Claude. We show that our REC models outperform state-of-the-art baselines on these benchmarks, and that the baselines often struggle to provide accurate citation, undermining their reliability. Specifically, our REC-70B model achieves Rank #1  [^5] As of Feb 15th, 2025 as a generative model on the RewardBench leaderboard under the model name TextEval-Llama3.1-70B. The model outperforms the second-ranked model in three dimensions: Chat Hard, Safety, Reasoning, which are key capabilities required to generate accurate explanation and citation outputs.
2 REC Model Citation Capabilities
In this section, we discuss the two main citation tasks, content quality citation and RAG citation, that the REC models support, with the understanding that the models can also be used for general purposes, such as question answering (QA) and summarization.
We illustrate the content quality citation task in Figure 1, where the autoevaluator performs content quality evaluation on another LLM’s generation according to a specified metric, and provides rating and explanation with citation for verifiability as an output. When not in an evaluation setting, our models can perform general RAG citation, where we cite from retrieved articles and tag the citation with an answer generated by another LLM, as illustrated in Figure 2.
Our models support various citation modes dependent on the citation tasks: post-fix citation (RAG only), inline citation (RAG only), post-fix citation with context snippet (for both content quality evaluation and RAG), and inline citation with context snippet (for both content quality evaluation and RAG). Output examples of different citation modes can be found in Appendices A and B.
These different modes were designed to accommodate different trade-offs between latency and granularity of citation. For instance, the post-fix citation mode would be the fastest as there is no need for our models to generate claims or snippets. It simply has to generate the reference to the cited context. On the other hand, the inline citation with context snippet mode has to place the citation inline with the generated response and also point to a snippet within the corresponding cited context. This mode is the most granular, but can increase latency as it needs to generate more output tokens.
@@ 3 Related Work
Providing a reference to the attributable source of information (Liu et al., 2023a; Gao et al., 2023) has been proposed as one of the approaches to mitigate hallucination and improve the factuality of LLM generation. Such work fall into the general RAG citation setting defined in our work, where the aim is to generate verifiable content through citation of provided articles. Initial efforts fine-tune LLMs using human-written examples (Nakano et al., 2021) or machine-generated examples verified by humans (Menick et al., 2022), but their privately maintained training data limits further research. With the advent of more capable LLMs (OpenAI, 2023; Jiang et al., 2023b), most existing work rely on zero-shot prompting or few-shot prompting to cite articles during generation (Kamalloo et al., 2023; Gao et al., 2023; Liu et al., 2023a), although the quality of their generated citation leaves much room for improvement (Malaviya et al., 2024). Other work utilize an additional natural language inference (NLI) model to add citation (Gao et al., 2022; Chen et al., 2023). To improve upon prior work, we propose a fine-tuned LLM that can generate better-grounded responses supported with citation of various modes, with the fine-tuning dataset released to the public to facilitate future research and no additional NLI models needed.
@@
Collecting human annotations to evaluate LLM is not only costly and time consuming, but hard to replicate (Ouyang et al., 2022; Zheng et al., 2023; Chiang and Lee, 2023), and as a result, LLMs become a natural automated proxy to evaluate LLM capabilities on various benchmarks (Bai et al., 2024; Bubeck et al., 2023; Chiang et al., 2023; Fu et al., 2023; Liu et al., 2023b; Wang et al., 2023a). Existing LLM autoevaluators often judge LLM outputs by expressing “preference” over outputs from a reference model (Dubois et al., 2024; Li et al., 2023b; Yuchen Lin et al., 2024), or by providing rating and explanation according to a user-defined metric as a direct assessment (Li et al., 2023a; Wang et al., 2023b, 2024b). Closer to our work, Jiang et al. (2024) and Xu et al. (2023) fine-tune LLMs to generate rating, explanation, and a detailed analysis to pinpoint errors in a response evaluated. Unlike prior work, our fine-tuned model is a more generalizable autoevaluator since it provides content quality citation for both good and bad evaluated responses, in addition to rating and explanation. This not only allows users to be able to diagnose where errors are from but provides supporting evidence when a generation is good, which are both important to establish trust in a model.
@@ 4 Data Collection
To fine-tune our general-purpose LLM autoevaluator, we meticulously collect a mixture of data encompassing a broad spectrum of LLM capabilities (Section 4.1), and we denote our collected dataset as REC-Data. Due to a lack of publicly available content quality citation datasets, we leverage synthetic data generation using an LLM to curate such data (Section 4.2) to facilitate the research community. We perform rule-based automatic quality check followed by a unified task formatting to post-process our data for instruction fine-tuning (Section 4.3).
@@ 4.1 Task Types
To enhance the explainability in the automatic evaluation of our LLM autoevaluator, while maintaining its general instruction-following capability, we gather datasets from a diverse range of task types (See detailed REC-Data distribution in Appendix C). These tasks represent essential capabilities that an advanced autoevaluator LLM should possess:
@@
  • Citation: Evaluate a response alongside the context, and provide citation for verifiable answer attribution.
@@ 4.2 Synthetic Data Generation
We leverage synthetic data generation for some of the tasks, including pointwise evaluation and citation. To ensure the quality of synthetic data, we leverage a powerful instruction-tuned model based on Llama-3.1-70B to generate both pointwise evaluation data and citation data for RAG and content quality.  [^6] Agreement between this model’s judgment and human labelers’ judgment is 86%, on par with the agreement between human labelers. For pointwise evaluation, the model rates every response on each of the 4 metrics listed in Section 1. Example prompts used to generate the synthetic data are provided in Appendix D.
@@ 4.3 Post-processing and Unified Task Format
Knowing that LLM-generated outputs are susceptible to hallucination, we further post-process the synthetic data to ensure the quality. Specifically, we filter synthetic LLM outputs that are either in wrong JSON output formats, or with invalid citation by checking whether the citation is extracted verbatim from the context. Our final REC-Data has in total around 140k datapoints, each containing a task prompt and a completion.
5 Model Training
Our general-purpose LLM autoevaluator is available in three sizes, REC-8B, REC-12B, and REC-70B. REC-12B is instruction fine-tuned from Mistral-Nemo [^7] https://mistral.ai/news/mistral-nemo, and REC-8B and REC-70B are instruction fine-tuned from Llama-3.1-8B and Llama-3.1-70B Team (2024), respectively. We trained REC-8B and REC-12B to assess the generalizability of the REC approach to different model architectures in similar model sizes. We adopted supervised fine-tuning (SFT) to optimize the models. To accommodate GPU memory constraints: we filtered examples where the combined length of a prompt and a response exceeded 6,144 tokens. We used Low-Rank Adaptation (LoRA) during fine-tuning Hu et al. (2021), with a rank of $r=256$ for REC-8B and REC-12B, and $r=64$ for REC-70B. We applied 4-bit quantization for REC-70B during initialization. All models were trained using a learning rate of $1\times 10^{-4}$, with data shuffled and for a single epoch. REC-8B and REC-12B were trained with a batch size of 2 and a gradient accumulation factor of 8, whereas for REC-70B we used a batch size of 1 with the same gradient accumulation factor. All models were trained on eight NVIDIA H100 GPUs, each with 80GB of memory. The total training time was approximately 4.5 hours for REC-8B and REC-12B, and 5.5 hours for REC-70B.
6 Evaluation
In this section, we evaluate our REC models on diverse benchmarks assessing LLMs’ (1) RAG citation capability: ALCE (Gao et al., 2023), and ExpertQA (Malaviya et al., 2024); (2) content quality citation capability: human evaluation on ABCD summarization (Chen et al., 2021); (3) general capabilities: RewardBench (Lambert et al., 2024), and LLM-AggreFact (Tang et al., 2024a); (4) cognitive bias: CoBBLEr (Koo et al., 2024), each of which is an important measure to a general-purpose LLM autoevaluator. We compare our results against 7 SOTA LLMs, including Misral-7B (Mistral-7B-Instruct-v0.2(Jiang et al., 2023a), Mistral-Nemo (Mistral-Nemo-Instruct-2407), Llama-3.1-70B (Grattafiori et al., 2024), Claude-3-Opus (claude-3-opus-20240229), GPT-3.5 (gpt-3.5-turbo), GPT-4 (gpt-4-turbo(OpenAI et al., 2024), and GPT-4o (gpt-4o).